### v1

# Graph Condensation: A Survey

Abstract
The burgeoning volume of graph data poses significant challenges in storage, transmission, and particularly the training of graph neural networks (GNNs). To address these challenges, graph condensation (GC) has emerged as an innovative solution. GC focuses on synthesizing a compact yet highly representative graph, on which GNNs can achieve performance comparable to trained on the large original graph. The notable efficacy of GC and its broad prospects have garnered significant attention and spurred extensive research. This survey paper provides an up-to-date and systematic overview of GC, organizing existing research into four categories aligned with critical GC evaluation criteria: effectiveness, generalization, fairness, and efficiency. To facilitate an in-depth and comprehensive understanding of GC, we examine various methods under each category and thoroughly discuss two essential components within GC: optimization strategies and condensed graph generation. Additionally, we introduce the applications of GC in a variety of fields, and highlight the present challenges and novel insights in GC, promoting advancements in future research.
1 Introduction
Graph data is extensively utilized across a diverse range of domains, owing to its capability to represent complex structural relationships among various entities in the real world Wu et al. (2022); Bo et al. (2023). Notable applications include, but are not limited to, social networks Li et al. (2018); Jung et al. (2012), chemical molecular structures Guo et al. (2022), transportation systems Rahmani et al. (2023), and recommender systems Zheng et al. (2023a); Yu et al. (2023a). However, the exponential growth of data volume in these applications makes the management and processing of large-scale graph data a complex and challenging task. A particularly demanding issue is the considerable computational load associated with training graph neural networks (GNNs) on large-scale graph datasets Hamilton et al. (2017). These challenges are more pronounced in scenarios that require training multiple GNNs, such as hyper-parameter optimization Ding et al. (2022) and neural architecture search Oloulade et al. (2021). This underscores the growing importance for developing more efficient and effective methodologies for processing large-scale graph data.
Figure 1: Taxonomy and evaluation criteria of graph condensation.
A direct solution to these challenges is graph reduction, where appropriate simplification of graphs can not only expedite graph algorithms but also improve the efficiency of storage, transformation and retrieval processes for diverse graph analysis tasks. Early explorations into graph reduction primarily focused on graph sparsification Spielman and Srivastava (2008) and coarsening Loukas and Vandergheynst (2018) techniques. These methods aim to reduce graph size by eliminating redundant edges and merging similar nodes. While effective in maintaining the essential characteristics of large graphs, these approaches have notable limitations. They predominantly rely on heuristic methods, such as the largest principal eigenvalues Loukas and Vandergheynst (2018) or pairwise distances Bravo Hermsdorff and Gunderson (2019), which can result in poor generalization across different GNNs architecture or downstream tasks. Consequently, the focus of research has shifted towards more sophisticated and adaptable methods for graph reduction. A leading development is graph condensation (GC) (a.k.a. graph distillation). GC Jin et al. (2022b) aims to synthesizes a small but highly informative graph dataset, encompassing both node attribute and topology structure, to effectively represent large original graph. Through task-driven optimization, models trained on these small condensed datasets can achieve performance comparable to those trained on the large original graphs. For instance, GCond Jin et al. (2022b) can condense the graph of Flickr dataset to just 0.1% of its original size, while some certain GNNs models trained on this small graph manage to retain 99.8% of the original test accuracy. Given the notable efficacy of GC, it has been soon applied to a variety of applications Gao et al. (2024); Liu et al. (2023b); Ding et al. (2022) and has sparked a proliferation of follow-up research Zheng et al. (2023c); Xu et al. (2023), further broadening the reach and deepening the impact of GC.
Although significant strides have been made to enhance the effectiveness of GC methods, as a data-centric approach, the evaluation of GC extends beyond mere effectiveness, motivating a broader exploration of additional criteria in assessing the performance of GC. One critical aspect is the condensed graph’s ability to generalize across multiple GNN architectures and various downstream tasks Jin et al. (2022b). A GC method with robust generalization is particularly beneficial in scenarios that demand training heterogeneous models or handling diverse tasks Yang et al. (2023a). In addition, it is essential for the condensed graph to maintain unbiased representations compared to the original graph, which is particularly crucial in fields sensitive to bias, such as human resources Bourmpoulias et al. (2023), criminal justice Lettieri et al. (2023), and financial services Wang et al. (2021). Beyond these considerations, the scalability of GC also emerges as a key factor in practical applications Gao et al. (2023b). An efficient GC method enables rapid adaption in life-long graph learning scenarios, effectively handling the continuous growth and dynamic changes of graph data. These criteria play a pivotal role in the evaluation of GC and simultaneously serve as guiding principles for research motivations. Thus, we categorize methods based on these criteria as shown in Figure 1, to provide a structured taxonomy for a comprehensive field survey. This taxonomy can help researchers in comprehensively understanding the diverse motivations driving GC methods, thereby facilitating a clearer path toward achieving their specific objectives.
While recent surveys on related topics including dataset distillation Geng et al. (2023); Yu et al. (2023b); Sachdeva and McAuley (2023); Lei and Tao (2024), data-centric graph learning Yang et al. (2023b); Zheng et al. (2023b); Shabani et al. (2024), as well as graph neural network acceleration Zhang et al. (2023) have touched upon GC, they summarize GC as a peripheral application or future direction, lacking systematic updates on the latest research and a holistic view. Moreover, existing dataset distillation surveys focus on detailed optimization strategies, missing high-level perspectives and design objectives analysis. This hinders researchers from understanding novel GC for varying contexts. Recognizing the rapid evolution and growing significance of GC, there is an urgent need for a systematic survey that encapsulates the latest advancements and motivations behind methodological designs. In summary, our contributions are as follows:
  • We introduces a systematic taxonomy for GC, categorizing existing methods into four distinct groups based on key criteria: effectiveness, generalization, fairness, and efficiency. This taxonomy not only structures existing knowledge, but captures the underlying motivations driving GC research.
  • We provide a comprehensive and up-to-date review of the latest advancements in GC. Additionally, we delve into two fundamental aspects: the optimization strategies and condensed graph generation, facilitating an in-depth understanding of GC techniques.
  • We discuss a variety of practical GC applications and shed light on current challenges and future directions in this field. This discussion aims to guide researchers who are seeking to advance GC further.
The structure of this survey is outlined as follows: Section 2 delves into the definition, design criteria, and taxonomy of GC. Following this, Section 3 discusses the optimization strategies employed in GC methods. Section 4 provides a thorough investigation of each category of GC, as aligned with our taxonomy. In Section 5, we extensively discuss the methods involved in condensed graph generation. Section 6 is dedicated to summarizing the various applications of GC. Finally, we present the challenges and future directions in this field in Section 7.
2 Definition and Taxonomy
2.1 Problem Definition
Given a graph dataset $\mathcal{T}=\left(\mathcal{V},\mathcal{E}\right)$ with $|\mathcal{V}|=N$ nodes and $|\mathcal{E}|=M$ edges, its adjacency matrix, node feature matrix and task specific label are denoted as ${\mathbf{A}}$, $\mathbf{X}$ and $\mathbf{Y}$, respectively [^2] We unify definitions for node-level and graph-level graph datasets.. Graph condensation aims to create a small condensed graph dataset $\mathcal{S}=\left(\mathcal{V}^{\prime},\mathcal{E}^{\prime}\right)$ with $|\mathcal{V}^{\prime}|=N^{\prime}$ and $N^{\prime}\ll N$. Its adjacency matrix, node feature matrix and task specific label are denoted as ${\mathbf{A}^{\prime}}$, $\mathbf{X}^{\prime}$ and $\mathbf{Y}^{\prime}$, respectively. To facilitate the connection between the original graph $\mathcal{T}$ and the condensed graph $\mathcal{S}$, a relay graph model $f_{\theta}\left(\cdot\right)$, parameterized by ${\theta}$, is employed in the optimization process for encoding both graphs as shown in Figure 2. Then, the graph condensation is formulated as:
$\mathcal{S}=\arg\min_{\mathcal{S}}\mathcal{L}_{cond}\left(f_{\theta}\left({\mathcal{T}},{\mathcal{S}}\right)\right),$ (1)
where $\mathcal{L}_{cond}$ is the optimization objective for graph condensation, which varies in format and will be elaborated upon in Section 3.
Figure 2: Graph condensation procedure. The original graph and condensed graph are encoded by relay model $f_{\theta}$ and the condensed graph is optimized according to $\mathcal{L}_{cond}$.
2.2 Criteria for Graph Condensation
While the accuracy of a specific GNN model, trained on the condensed graph and evaluated with the original graph, is considered the most important metric for evaluating a GC method, there are other criteria that merit consideration. Specifically, we summarize the key criteria of GC as follows:
  • Effectiveness assesses the accuracy of models trained on condensed graphs at various compression rates, using a specified GNN architecture Jin et al. (2022b). It measures the extent to which the condensed graph preserves the essential characteristics of the original graph.
  • Generalization measures the overall accuracy of different GNN architectures trained on the condensed graph across various downstream tasks Yang et al. (2023a). It reflects the condensed graph’s adaptability and utility in diverse contexts.
  • Fairness evaluates the disparity in fairness between models trained on condensed graphs and those trained on original graphs Feng et al. (2023), aiming to prevent bias amplification in the condensed graph and ensure equitable model performance.
  • Efficiency assesses the time required to learn a condensed graph Jin et al. (2022a), focusing on the speed of condensed graph generation and the method’s feasibility for time-sensitive scenarios.
Given these criterion, an ideal GC method should generate a high-quality condensed graph dataset, enabling models with different GNN architectures to achieve results comparable and unbiased to those obtained on the original graph on a wide range of downstream tasks.
2.3 Taxonomy
The above criteria not only serve as the indicators of GC’s utility, but also act as the objectives that steer extensive research efforts in this field. To streamline the understanding of GC, we class GC methods into four categories which align with the above criteria: effective GC, generalized GC, fair GC, and efficient GC, as shown in Figure 1.
In addition to our proposed taxonomy, we also provide a detailed summary of GC methods, highlighting their distinct characteristics, in Table 1. This table offers a comprehensive comparison across several dimensions, including motivation, optimization strategy, condensed graph generation method, and downstream tasks.
Table 1: Summary of existing graph condensation research, organized chronologically by publication date. In optimization strategy: “GM” represents gradient matching, “DM” refers to distribution matching, “TM” is short for trajectory matching and “KRR” is kernel ridge regression. In relay model: “GNN” is graph neural network, “GNTK” refers to graph neural tangent kernel and “SD” is short for spectral decomposition. In downstream task: “NC” represents node classification, “GC” refers to graph classification, “AD” is anomaly detection and “LP” is short for link prediction.
Method Motivation Input Optimization strategy Output Downstream task Enhanced module
Pre-processing Objective Relay model Bi-level Structure construction Sparsification
GCond Jin et al. (2022b) Effectiveness GM GNN Generative model Threshold NC N/A
DosCond Jin et al. (2022a) Efficiency GM GNN Parameterization Gumbel softmax GC Optimization
GCDM Liu et al. (2022) Efficiency DM GNN Generative model Threshold NC Objective
HCDC Ding et al. (2022) Application GM GNN Generative model Threshold NC Condensed graph
MSGC Gao and Wu (2023) Effectiveness GM GNN Pre-defined structure N/A NC Condensed graph
SGDD Yang et al. (2023a) Generalization GM GNN Generative model Regularization NC, AD, LP Objective
SFGC Zheng et al. (2023c) Effectiveness TM GNN Pre-defined structure N/A NC Objective
FGD Feng et al. (2023) Fairness GM GNN Generative model Threshold NC Objective
GCARe Mao et al. (2023) Fairness GM GNN Generative model Threshold NC Relay model
MCond Gao et al. (2024) Application GM GNN Generative model Threshold NC Objective
KIDD Xu et al. (2023) Efficiency KRR GNTK Parameterization Gumbel softmax GC Relay model
CaT Liu et al. (2023b) Application DM GNN Pre-defined structure N/A NC Optimization
CTRL Anonymous (2023) Effectiveness GM GNN Generative model Threshold NC, GC Objective
FedGKD Pan et al. (2023) Application N/A GNN Generative model Gumbel softmax NC Condensed graph
GCEM Liu et al. (2023a) Generalization DM SD Parameterization N/A NC Relay model
GC-SNTK Wang et al. (2023) Efficiency KRR GNTK Parameterization N/A NC Relay model
GroC Li et al. (2023) Effectiveness GM GNN Generative model Threshold NC, GC Objective
PUMA Liu et al. (2023c) Application DM GNN Pre-defined structure N/A NC Objective
2.3.1 Effective Graph Condensation
The primary objective of GC focuses on optimizing the task-specific quality of the condensed graph. As illustrated in Figure 3, the condensed graph produced by effective GC methods enables GNNs to achieve performance that closely approximates those trained on the original graph, especially in targeted downstream tasks. Consequently, advanced effective GC significantly enhances the feasibility of using condensed graphs as substitutes. To systematically study these methods, we further divide this category into two distinct classes: optimization-driven methods and module-enhanced methods, providing a clearer and more detailed exploration of effective GC methodology.
Figure 3: Effective graph condensation (GC). The condensed graph enables GNNs to be trained with performance comparable to GNNs trained on the original graph.
2.3.2 Generalized Graph Condensation
The objective of GC extends beyond merely training specific models for designated tasks. It also involves enabling the application of various models across a spectrum of tasks, as depicted in Figure 4. A condensed graph with robust generalization capabilities avoids repeated generation of task-specific or model-specific condensed graphs, thereby significantly enhancing its utility in various applications, such as neural architecture search, hyper-parameter optimization, and multi-task learning. The primary challenge of generalization is to minimize the loss of information in both the structural and feature representations during condensation. Consequently, we study existing generalized GC from two primary aspects: advanced relay model for better information preservation and regularization technique to compensate information loss during condensation.
Figure 4: Generalized graph condensation (GC). The condensed graph facilitates the training of various GNN architectures and supports various downstream tasks.
2.3.3 Fair Graph Condensation
Balancing information during the condensation process represents a pivotal challenge in the domain of GC. The intensive compression of GC often leads to an amplification of data biases within the condensed graphs. Consequently, GNNs trained on such condensed graphs exhibiting more pronounced fairness issues compared to those trained on original graphs. In response to this, fair GC methods constrain the condensation procedure to mitigate bias in the condensed graph as shown in Figure 5. The approaches to improving fairness in GC encompass two key ideas. The first idea is to integrate the fairness-enhancing regularization into the GC objective function, effectively guiding the optimization process toward fairer outcomes. The second idea is the development of novel modules within the relay model, promoting unbiased representations in the condensed graphs.
Figure 5: Fair graph condensation (GC) ensures that GNNs trained on condensed graphs produce results unbiased relative to those from GNNs trained on the original graph.
2.3.4 Efficient Graph Condensation
An efficient condensing procedure is essential for rapid adaptation in life-long graph learning scenarios. However, GC often suffers from sophisticated optimization process and slow convergence, resulting in a time-consuming condensing procedure. This particularly deteriorates in life-long learning contexts, where regular updates to the condensed graph dataset are a common requirement, as depicted in Figure 6. To address these challenges, efficient GC methods have been developed to accelerate the condensation process across all components within GC . To thoroughly explore the mechanisms underlying efficient GC, we break down the conventional GC procedure into three stages: graph encoding, optimization, and graph generation. These components align with the input, processing, and output stages of GC, respectively. By enhancing the efficiency of each stage, efficient GC methods significantly reduce the time required for the entire GC process.
Figure 6: Efficient graph condensation (GC) contributes to accelerating the condensation procedure.
3 Optimization Strategies
As we delve deeper into the intricacies of GC, it is crucial to examine the optimization strategies that form a core module within the GC framework. We hereby introduce the various optimization strategies employed in GC, offering a high-level perspective on the condensation procedure. Our discussion commences with the foundational principle of optimization objective (Eq. (3)), then move onto detailed optimization strategies (Eq. (4)-(8)).
The foundational principle of the optimization strategies posits that the performance of a model trained on the condensed data should align with that of a model trained on the original data. To this end, we initially define the loss for data $\mathcal{T}$ and $\mathcal{S}$ in relation to the parameter ${\theta}$ of the relay model as:
$\begin{split}&\mathcal{L}^{\mathcal{T}}\left(\theta\right)=\ell\left(f_{\theta}\left({\mathcal{T}}\right),\mathbf{Y}\right),\\ &\mathcal{L}^{\mathcal{S}}\left(\theta\right)=\ell\left(f_{\theta}\left({\mathcal{S}}\right),\mathbf{Y}^{\prime}\right),\end{split}$ (2)
where $\ell$ is the task-specific objective such as cross-entropy. Then the objective of GC can be formulated as the following bi-level problem:
$\min_{\mathcal{S}}\mathcal{L}^{\mathcal{T}}\left(\theta^{\mathcal{S}}\right)\>\>\text{s.t.}\>\>\theta^{\mathcal{S}}=\arg\min_{\mathcal{\theta}}\mathcal{L}^{\mathcal{S}}\left(\theta\right),$ (3)
where the inner loop optimizes the parameter $\theta$ of relay model on the condensed data until convergence, and the outer loop subsequently optimizes the condensed data $\mathcal{S}$. Addressing the above objective necessitates the incorporation of the task-specific objective $\ell$ and entails resolving a nested loop optimization, which includes unrolling the entire training trajectory of the relay model optimization. This process can be prohibitively expensive. To alleviate the complexity of the optimization process, various approximation methods have been proposed, including gradient matching, trajectory matching, kernel ridge regression and distribution matching.
Gradient matching. Gradient matching is initially proposed by Zhao et al. (2020) and formulates the condensation objective in Eq. (3) as matching the optimized parameters of the models trained on two datasets:
$\min_{\mathcal{S}}\mathbb{E}_{\theta_{0}\sim{\Theta}}[\mathcal{D}\left({\theta}^{\mathcal{T}},{\theta}^{\mathcal{S}}\right)],$ (4)
where $\theta_{0}$ is the initialization of ${\theta}^{\mathcal{T}}$ and ${\theta}^{\mathcal{S}}$. ${\Theta}$ is a specific distribution for relay model initialization. The expectation on $\theta_{0}$ aims to improve the robustness of ${\mathcal{S}}$ to different parameter initialization Lei and Tao (2024). $\mathcal{D}\left(\cdot,\cdot\right)$ is a function measuring the distance between the gradients w.r.t. ${\theta}^{\mathcal{T}}$ and ${\theta}^{\mathcal{S}}$. The bi-level objective in Eq. (3) is approximated by matching the model gradients at each training step $t$. In this way, the training trajectory on condensed data can mimic that on the original data, i.e., the models trained on these two datasets converge to the similar solutions. The optimization objective $\mathcal{L}_{cond}$ in Eq. (1) is defined as:
$\begin{split}\mathcal{L}_{cond}=\mathbb{E}_{\theta_{0}\sim{\Theta}}&\left[\sum_{t=1}^{T}\mathcal{D}\left(\nabla_{\theta}\mathcal{L}^{\mathcal{T}}\left({\theta_{t}}\right),\nabla_{\theta}\mathcal{L}^{\mathcal{S}}\left({\theta_{t}}\right)\right)\right]\,\,\\ &\text{s.t.}\,\,{\theta}_{t+1}=\operatorname{opt}\left(\mathcal{L}^{\mathcal{S}}\left({\theta_{t}}\right)\right),\end{split}$ (5)
where $\operatorname{opt}\left(\cdot\right)$ is the model parameter optimizer and the parameter of relay model is updated only on $\mathcal{S}$.
Trajectory matching. Gradient matching primarily aligns single-step gradients, yet it may accumulate errors when the relay model is iteratively updated using condensed data across multiple steps. To mitigate this problem, Cazenavette et al. (2022) propose a multi-step matching approach to approximate Eq. (4) and known as trajectory matching.
Trajectory matching trains two separate relay models on condensed data and original data, respectively. These two relay models are trained for distinct steps starting from a same initialization $\theta^{\prime}_{t}$, which is sampled from the original data training checkpoints ${\Theta^{\prime}}$. Then, trajectory matching aims to minimize the discrepancy between the final points of these two training trajectories:
$\begin{split}&\mathcal{L}_{cond}=\mathbb{E}_{\theta^{\prime}_{t}\sim{\Theta^{\prime}}}\left[\mathcal{D}\left({\theta}^{\mathcal{T}}_{t+k},{\theta}^{\mathcal{S}}_{t+l}\right)\right]\\ &\text{s.t.}\,\,{\theta}^{\mathcal{T}}_{t+1}=\operatorname{opt}\left(\mathcal{L}^{\mathcal{T}}\left({\theta^{\mathcal{T}}_{t}}\right)\right),\\ &\,\,\,\,\,\,\,\,\,{\theta}^{\mathcal{S}}_{t+1}=\operatorname{opt}\left(\mathcal{L}^{\mathcal{S}}\left({\theta^{\mathcal{S}}_{t}}\right)\right),\end{split}$ (6)
where $k$ and $l$ are hyper-parameters controlling the update steps. ${\theta}^{\mathcal{S}}_{t}$ and ${\theta}^{\mathcal{T}}_{t}$ denote the parameters of relay models trained on $\mathcal{T}$ and $\mathcal{S}$ at step $t$ respectively. Although trajectory matching introduces a tri-level optimization process, the training phase on the original data can be pre-processed. Consequently, the online condensation procedure remains a bi-level optimization process.
Kernel ridge regression. Gradient matching and trajectory matching approximately optimize the objective Eq. (4) by matching one-step gradient or multi-step trajectory. However, these inexact solutions can hardly lead to the optimal condensed graph. To bridge this gap, Nguyen et al. (2020) transform the original task into a regression problem by substituting the neural network with kernel ridge regression (KRR). This approach enables the direct resolution of the constraint in Eq. (3) using a closed-form solution. Consequently, the complex nested optimization is simplified into a first-order process by considering the condensed data as the support set and the original data as the target set. The optimization problem described in Eq. (3) is reformulated into a regression task:
$\mathcal{L}_{cond}=\frac{1}{2}\lVert{\bf Y}-K_{{\mathcal{T}}{\mathcal{S}}}\left(K_{{\mathcal{S}}{\mathcal{S}}}+\lambda I\right)^{-1}{\bf Y}^{\prime}\rVert^{2},$ (7)
where $K_{{\mathcal{T}}{\mathcal{S}}}$ is the kernel matrix of ${\mathcal{T}}$ and ${\mathcal{S}}$, and $K_{{\mathcal{S}}{\mathcal{S}}}$ is the kernel matrix of ${\mathcal{S}}$. It is important to note that matrix inversion is solely dependent on the number of condensed samples, making it computationally efficient.
Distribution matching. All aforementioned optimization strategies stem from the objective of aligning model performance as outlined in Eq. (3), inevitably coupling the relay model with the optimization process. To avoid this issue, Zhao and Bilen (2023) propose an alternative optimization approach known as distribution matching, which focuses on matching the feature distributions of condensed and original data. The objective of distribution matching is to generate condensed data with a feature distribution closely approximating that of the original data. Accordingly, the optimization objective of distribution matching is formulated as:
$\mathcal{L}_{cond}=\mathbb{E}_{\theta_{0}\sim{\Theta}}\left[\mathcal{D}\left(f_{\theta}\left({\mathcal{T}}\right),f_{\theta}\left({\mathcal{S}}\right)\right)\right].$ (8)
To address the inconsistency in sample quantities between original and condensed dataset, the distance function $\mathcal{D}\left(\cdot,\cdot\right)$ is specifically designed for class-wise comparison. Moreover, distribution matching eliminates the requirement for the task-specific objective in Eq. (2), thereby enhancing the efficiency of the condensation process.
4 Graph Condensation Methods
In this section, we dive into the specifics of GC and conduct a systematic analysis of four distinct GC methodologies: effective GC, generalized GC, fair GC, and efficient GC, exploring the techniques employed to achieve their respective objectives.
4.1 Effective Graph Condensation
As mentioned in Section 2, we further divide effective GC methods into two categories according to the enhanced components within GC framework, including optimization-driven methods and module-enhanced methods.
4.1.1 Optimization-Driven Methods
Designing more advanced optimization strategies is often considered the critical driving force to achieve effective GC, thus steering many studies in this field. The first GC method Jin et al. (2022b) employs the gradient matching as the optimization strategy. However, the limitations associated with single-step gradient alignment constrain the overall performance of GC. Therefore, the follow-up studies are proposed to explore and enhance the optimization process. For example, CTRL Anonymous (2023) argues that the distance measurement in gradient matching is coarse, which relies on cosine similarity and solely focuses on the directional aspect of gradients. This fails to adequately capture the vector characteristics of gradients, thereby introducing biases in the gradient matching process. Therefore, CTRL proposes to refine the matching objective by adding the magnitudes distance. Furthermore, CTRL empirical discovers that gradient magnitude matching not only explicitly accounts for the vector properties of gradients, but also implicitly benefits the alignment of frequency distributions between condensed and original graphs. This provides another compelling rationale for the performance improvement.
Despite the improvements from fine-grained gradient matching, CTRL is still focuses on short-range matching. This leads to the short-sight issue and fails to capture the holistic GNN learning behaviors, potentially compromising the quality of condensed graphs. To address this limitation, SFGC Zheng et al. (2023c) introduces trajectory matching in GC and proposes to align the long-term GNN learning behaviors between the original graph and condensed graph. As a result, it enables a more comprehensive and efficient knowledge transfer to the condensed graph.
Beyond aligning gradients or trajectories during model training, kernel ridge regression is employed for precise solutions in bi-level optimization objectives. For instance, GC-SNTK Wang et al. (2023) discards the GNN and employ the kernel ridge regression, incorporating graph neural tangent kernels Du et al. (2019) as the relay model. This modification not only facilitates the first-order optimization of the condensed graph, but also eliminates the requirement for multiple initializations of relay GNN, resulting in a more stable condensation procedure. KIDD Xu et al. (2023) employs kernel ridge regression for graph classification tasks and further improves the method by removing the non-linear activation in graph neural tangent kernels. This approach not only simplifies the matrix multiplications in the iterative updates of condensed graphs but also leads to substantial improvements in model accuracy compared to the gradient matching method.
4.1.2 Module-Enhanced Methods
Optimization-driven methods enhance the effectiveness of GC by emphasizing the optimization procedure. In contrast, other studies improve the performance by augmenting the other modules with in GC framework.
As a representative method in this category, MSGC Gao and Wu (2023) improves the GC from graph structure construction. Unlike many GC approaches which synthesize a single dense adjacency matrix, MSGC instead pre-defines multiple small-scale sparse graphs. These pre-defined structures allow each condensed node encompasses distinct neighborhoods. The variation in structure enables GNNs to capture a broader information, leading to diverse node embeddings and enhanced performance. PUMA Liu et al. (2023c) emphasises the significance of label quantity in GC and improves the condensation quality by incorporating pseudo labels from the original graph. The inclusion of additional supervision signals enhances class representation and ensures more accurate information preservation in the condensed graph. Moreover, GroC Li et al. (2023) designs a perturbation module in the process of generating condensed graphs. This module maximizes the distance between the condensed graph and the original graph by selectively perturbing the condensed graph that are insufficiently informative. By utilizing adversarial training technique, these perturbations are capable of exploring the neighborhood areas within the parameter space of the condensed graph and significantly enhancing the gradient matching process.
4.2 Generalized Graph Condensation
The essence of generalization lies in preserving crucial task-related information while mitigating bias injection into the condensed graph. Nevertheless, some GNN relay models, employed for encoding both structure and attribute information, inevitably result in a loss of structural information and create an entanglement between the condensed graph and the relay model. These issues significantly affect the generalization ability to unseen models and tasks. To explore the cause, GCEM Liu et al. (2023a) conducts an investigation across various GNN architectures, including different spatial and spectral GNNs. They find that the performance gap increases when the relay model and downstream model have different architectures. This is due to that distinct spectral preferences of GNNs have varying impacts on the significance of different graph eigenvectors Bo et al. (2023); Zhu et al. (2021), consequently leading to alterations in node features. For instance, if a GNN functions primarily as a low-pass filter, it will prioritize low-frequency information in node features, overlooking the full spectrum. The loss of high-frequency information will compromise the high-pass filter GNNs’ performance.
To address this issue, generalized GC methods are developed from the spectral perspective. For example, GCEM Liu et al. (2023a) circumvents the conventional relay GNN and directly generates the eigenbasis for the condensed graph, thereby eliminating the spectrum bias inherent in relay GNNs. To address the inconsistencies in eigenvector quantity and dimension, GCEM aligns the condensed eigenvectors with the selected eigenvectors of the original graph and matches the class features in the subspace defined by these eigenvectors.
Apart from leveraging the advanced relay model, SGDD Yang et al. (2023a) employs the Laplacian energy distribution (LED) to analyze the structural properties of the graph, quantifying the spectral shift between condensed and original graphs. A substantial LED shift suggests that crucial structural information from the original graph is lost, diminishing the generalization capabilities of the condensed graph. Therefore, SGDD leverages the LED as a structural shift metric and designs an additional regularization to enhance the structure of the condensed graph. This regularization aligns the LED between the original graph and the condensed graph, facilitating the broadcast of structural information. By preserving crucial structural information, SGDD effectively reduces the performance discrepancies observed across different GNN architectures and various tasks, e.g. node classification, anomaly detection and link prediction.
4.3 Fair Graph Condensation
GC can effectively capture the essential task-specific characteristics of the original graph by exclusively optimizing for task performance. However, this task-centric optimization approach may introduce fairness concerns, primarily attributed to the indiscriminate transmission of information into the condensed graph. Especially in scenarios with intensive compression rates, GC tends to amplify biases as it prioritizes the preservation of predominant and representative information. Consequently, GNNs trained on condensed graphs exhibit more severe fairness issues than those trained on original graphs Feng et al. (2023); Mao et al. (2023), which challenges the applicability of GC in high-stake applications Mehrabi et al. (2021); Feng et al. (2023).
To mitigate bias in condensed graphs, FGD Feng et al. (2023) initially investigates the relationship between the feature spaces of original graph and condensed graph, and constructs an estimator for sensitive attributes. Following this, the variance of the estimated sensitive group membership is employed as the bias measurement and integrated as the regularization to facilitate the unbiased condensed graph. Moreover, GCARe Mao et al. (2023) directly constrains the condensation process with adversarial training. It additionally incorporates a discriminator into the relay model, which is designed to predict the sensitive groups of node representations. The relay model operates as a generator, generating the fair data to deceive the discriminator. Consequently, the condensation process is formulated as a min-max optimization problem, striking a balance between data generation and debiasing efforts.
4.4 Efficient Graph Condensation
The GC process requires significant computational resources. To figure out the root causes of heavy computation, we break down the GC process into three fundamental stages, including: efficient graph encoding, efficient optimization and efficient graph generation, and examine acceleration methods for each component.
4.4.1 Efficient Graph Encoding
In GC, the graph encoding process is executed by the relay model. Generally, relay models used in GC methods adopt either the message-passing framework or the graph neural tangent kernel framework. All these frameworks stack multiple propagation and transformation operations to aggregate and update information according to the graph structure. However, as multiple propagation processes are executed, the number of aggregated neighboring nodes increases exponentially, leading to significant computational demands.
To mitigate the expensive computation resulting from propagation, some GC methods borrow the idea from SGC Wu et al. (2019) and remove the non-linear transformation layers between aggregation layers in the relay model. For example, Gcond Jin et al. (2022b) and PUMA Liu et al. (2023c) utilize SGC to encode the original graph. KIDD Xu et al. (2023) simplifies the graph neural tangent kernel by removing non-linear activation at certain layers. By this means, the propagation of these methods can be pre-computed in the pre-processing stage and only needs to be executed once. Instead of performing propagation during each training epoch, the online time complexity of GC is significantly reduced than utilizing other relay models.
4.4.2 Efficient Optimization
In GC, the optimization procedure often requires numerous iterations to update the relay model before optimizing the condensed graph, leading to considerable computation.
To mitigate this problem, DosCond Jin et al. (2022a) proposes the one-step gradient matching strategy which discards the relay model update and only performs gradient matching for one single step. The theoretical analysis posits that minimizing the one-step matching loss serves a dual purpose: it not only directs the learning process of condensed graphs towards lower losses on original graphs but also delineates the trajectory for updating the condensed data. Moreover, one-step matching eliminates the need for hyper-parameter tuning, such as the number of iterations of bi-level optimization, further simplifying the condensation process.
Instead of matching gradients, distribution matching methods change the optimization strategy towards approximating original graph in the feature space. This approach leads to a shift in the role of the relay model within GC. It transitions from aligning task performances to transforming both original and condensed graphs into a unified space for distribution alignment. Consequently, distribution matching obviates the need to compute the gradients and trajectories w.r.t. parameters of the relay model in each training iteration, leading to a substantial reduction in computational complexity. Specifically, GCDM Liu et al. (2022), CaT Liu et al. (2023b), PUMA Liu et al. (2023c) and GCEM Liu et al. (2023a) compute node distributions for both original and condensed graphs in the same feature space and employ the maximum mean discrepancy Gretton et al. (2012) between each pair of class distributions as the optimization objective.
Another research line to avoid the bi-level optimization is kernel ridge regression-based methods Xu et al. (2023); Wang et al. (2023). These methods reformulate the classification task as a kernel ridge regression task and incorporate the graph neural tangent kernel within kernel ridge regression to encode the original graph. Kernel ridge regression provides an efficient closed-form solution, dispensing with the necessity for iterative computations in updating the relay model and resulting in significantly faster convergence speeds.
4.4.3 Efficient Graph Generation
Condensed graph generation targets two components: the feature matrix and the adjacency matrix. Modeling these matrices typically requires two distinct optimization processes. For example, GCond Jin et al. (2022b) models the condensed structure as a function of the condensed nodes, requiring the optimization of an additional adjacency matrix generator. To simplify the optimization procedure, several studies Jin et al. (2022b); Zheng et al. (2023c); Liu et al. (2022, 2023b, 2023c) discard the adjacency matrix and instead use an identity matrix to represent condensed structure. These structure-free GC methods implicitly encode the original graph’s topological structure into discriminative node attributes via relay GNNs and these enriched embeddings are informative to train different downstream GNNs.
5 Condensed Graph Generation
Different from traditional graph reduction methods that trim the original graph, GC focuses on creating an entirely new condensed graph. An effective graph generation method not only accelerates optimization convergence but also enhances overall performance. In this section, we discuss the condensed graph generation methods from two aspects: attribute initialization and structure construction, as shown in Fig. 7.
Figure 7: Condensed graph generation methods.
5.1 Attribute Initialization
Diverse initializations of the attribute matrix $\mathbf{X}^{\prime}$ can lead to varying convergence performance. To expedite the optimization process for the attribute matrix, researchers have investigated a range of initialization methods as substitutes for the conventional random initialization. For instance, GCond Jin et al. (2022b) initiates attribute matrix by randomly selecting original nodes with corresponding labels. SFGC Zheng et al. (2023c) employs the coreset method to select the most representative nodes for initialization. Besides the node selection, Mcond Gao et al. (2024) and CTRL Anonymous (2023) demonstrate that employing the synthetic embeddings can further enhance optimization performance. Accordingly, these methods first cluster original graph nodes based on the pre-defined compression rate, and then utilize the cluster centroids as the initialization of attribute matrix.
5.2 Structure Construction
The adjacency matrix $\mathbf{A}^{\prime}$ of the condensed graph reflects relationships among condensed nodes and preserves the topological information of the original graph. The construction method of adjacency matrix affects its characteristics, such as homophily, spectral properties, and sparsity, which in turn impact overall performance. Current studies employ three kinds of strategies for generating adjacency matrices: generative model, parameterization, and predefined structure.
Generative model. A straightforward generation method involves employing the generative model to construct adjacency matrix explicitly. For example, SGDD Yang et al. (2023a) utilizes noise, along with the attributes and labels of the condensed graph, as inputs to the generative model for creating the adjacency matrix. Although this method offers greater flexibility in controlling the generated structures, it suffers from the substantial modeling complexity: the number of parameters in $\mathbf{A}^{\prime}$ grows quadratically with the increase of condensed nodes.
Parameterization. To mitigate modeling complexity of graph structure, several GC methods propose to parameterize the adjacency matrix. GCond Jin et al. (2022b) constructs the adjacency matrix based on homophily and models the adjacency matrix by node similarities. This parameterization strategy is designed to investigate the correlations between pairs of condensed nodes. KIDD Xu et al. (2023) directly optimize the low-rank decomposed matrices of the adjacency matrix to speed up the convergence. To preserve the spectral characteristics, GCEM Liu et al. (2023a) decomposes the adjacency matrix into eigenbasis and constructs the condensed structure using the eigenvectors matched with the most significant eigenvectors of the original graph.
Pre-defined structure. To further simplify the modeling of condensed graph, several studies leverage predefined structure. For instance, MSGC Gao and Wu (2023) pre-defines multiple sparse adjacency matrices for condensed nodes. By this means, MSGC reduces the complexity of the adjacency matrix while capturing a more diverse range of structural information compared to the single dense matrix. Moreover, several GC methods have incorporated the structure-free strategy Jin et al. (2022b); Zheng et al. (2023c); Liu et al. (2022, 2023b, 2023c) by employing a fixed identity matrix as the adjacency matrix. This strategy leads to a significant simplification of the computational requirements for adjacency matrix.
5.2.1 Condensed Graph Sparsification
The structure construction methods typically yield dense adjacency matrices with continuous values. These dense matrices often contain small but storage-intensive values that have minimal impact on GNN aggregation performance. Moreover, certain application scenarios, such as molecular graphs, require the binary and discrete adjacency matrices for downstream tasks. Hence, it becomes essential to sparsify the generated adjacency matrices, and numerous sparsification methods have been proposed to tackle these challenges. For instance, GCond Jin et al. (2022b) utilizes the threshold to remove the entries with small values and justifies that suitable choices of threshold do not degrade performance but increase the scalability of condensed graph. SGDD Yang et al. (2023a) employs the L2 norm of the adjacency matrix as a regularization term for sparsity control. To construct binary and discrete adjacency matrices, DosCond Jin et al. (2022a) and KIDD Xu et al. (2023) model each pair of nodes as an independent Bernoulli variable. Then, the Gumbel-Max reparametrization trick Maddison et al. (2016) is utilized to address the non-differentiability challenges in the optimization procedure.
6 Applications
Due to the efficiency in training GNN models and capacity to conserve storage costs, GC contains vast application prospects and is increasingly being utilized across a variety of fields. In this section, we enumerate the diverse applications of GC, including hyperparameter/neural architecture search, graph continual learning, federated learning, inference acceleration and recommender systems.
6.1 Hyper-Parameter/Neural Architecture Search
The hyper-parameter/neural architecture search Gao et al. (2021) involves systematic experimentation with a range of hyper-parameters or architectures to identify the optimal combination for downstream task. This process typically requires the training of multiple models with varying hyper-parameters/architectures on the same dataset, incurring substantial computational costs. To mitigate the computational burden, HCDC Ding et al. (2022) introduces GC in this problem and additionally generates the synthetic validation data by matching the hyper-parameter gradients. This ensures that the performance ranking of hyper-parameters/architectures on condensed datasets is consistent with that on the original datasets. By integrating GC, HCDC significantly reduces the time required for the search process.
6.2 Graph Continual Learning
Graph continual learning Febrinanto et al. (2023) is the approach that focuses on the continuous acquisition of knowledge from the dynamic and evolving graph data. In this realm, a critical challenge is the catastrophic forgetting problem, where a model’s performance on previously learned data deteriorates significantly when trained on new data. To mitigate this issue, CaT Liu et al. (2023b) and PUMA Liu et al. (2023c) incorporate GC into the process of graph continual learning, leveraging the condensed graphs to preserve historical information within a memory bank. To maintain a balance between historical and newly incoming data, CaT strategically trains the model exclusively on the condensed graph. This method not only addresses the challenge of data imbalance in graph continual learning but also significantly enhances learning efficiency.
6.3 Federated Learning
Federated learning Fu et al. (2022) is a distributed model training framework that safeguards private data, involving multiple clients with exclusive data and a central server. In this setup, clients train local models with their unique data and then transmit the parameters to the server, which aggregates and redistributes the updated knowledge. A key challenge in federated learning is the non-i.i.d nature of client data, leading to biased local models and slow convergence, increasing the communication load between the server and clients. To address this, FedGKD Pan et al. (2023) integrates GC at the client level to efficiently extract the local task features, enhancing the similarity computation and ensuring better incorporation of local information. Combined with global aggregation, FedGKD significantly improves collaboration results in federated learning environments.
6.4 Inference Acceleration
In practical graph systems, new present nodes will be connected with existing nodes in the original graph, then the message passing process is executed to aggregate information and generate representations for inference. However, the extensive size of large original graphs poses a considerable challenge for real-time inference Gao et al. (2023b). This difficulty primarily stems from the “neighbor explosion” problem, where the number of aggregated nodes exponentially increases with the propagation depth. To reduce the high computational demands and latency inherent in GNNs, MCond Gao et al. (2024) adopts GC and proposes to learn a mapping from original nodes to condensed nodes. This approach allows direct information propagation on the small condensed graph, resulting in a remarkable acceleration of the inference process.
6.5 Recommender Systems
Recommender systems are widely used in various industries such as e-commerce and social media, aiming to suggest relevant items or content to users. In a typical recommender system, the user-item interactions can be represented as a bipartite graph. As the number of users and items grows, this graph becomes increasingly large and complex, posing significant computational challenges. To handle this issue, Distill-CF Sachdeva et al. (2022) applies GC to recommender systems. It adopts a multi-step Gumbel sampling trick to condense the original user base into a limited number of representative users, which are then iteratively updated through distilling knowledge from the original graph. By replacing the original user-item graph with the synthesized small graph for training and inference, this method leads to more scalable and efficient recommender systems.
7 Challenges and Future Directions
Although current GC methodologies have achieved great success across various applications, GC still grapples with inherent limitations and unresolved challenges. In this section, we summarize the current challenges and identify several future research directions worth exploration.
7.1 Condensation for Diverse Graphs
As an innovative approach in graph reduction, current GC methods have predominantly focused on simple attribute graphs. However, the graph data in practical applications is far more complex, showcasing a wide spectrum of characteristics including heterophilic graphs Zheng et al. (2022), heterogeneous graphs Gao et al. (2023a), digraphs Bang-Jensen and Gutin (2008), and dynamic graphs Kazemi et al. (2020), etc. Each type of graph brings forth its own unique features, necessitating a deep understanding of their structures and a requirement for diverse GC strategies. Such a comprehensive grasp is not only critical for effectively managing the increasing volumes of data but also plays a key role in widening the horizons of their applications.
7.2 Task-Agnostic Graph Condensation
The optimization strategies currently employed in GC are specifically designed for the classification task, heavily relying on class labels. However, the entanglement of the GC optimization process with task-related labels can potentially hinder performance in other downstream tasks. Therefore, there is a noteworthy potential to develop generalized GC methods for diverse downstream tasks, including link prediction Kumar et al. (2020), anomaly detection Akoglu et al. (2015), graph reconstruction Xia et al. (2021), and graph generation Zhu et al. (2022), etc. Furthermore, considering the extensive time and resources required for data annotation, the development of self-supervised GC can significantly advance and expand the adaptability of GC techniques.
7.3 Efficient Evaluation for Condensed Graphs
In the realm of GC, the primary objective is to create high-quality condensed graphs. However, a systematic and well-established methodology for assessing the condensed graph quality is notably absent in the field Zheng et al. (2023c). This gap is particularly critical given the complex optimization involved in GC. The assessment of the condensed graph quality is essential not only for monitoring the condensation progress but also for potentially streamlining the optimization through early-stop strategies. Currently, GC practices leverage the downstream GNNs performance as the indicator for assessment. However, this method adds considerable computational overhead to the GC process. Moreover, the concept of quality in condensed graphs is inherently complex, varying significantly across different applications and necessitating a multifaceted approach to evaluation. Considering these challenges, there is a pressing requirement to develop the comprehensive and efficient evaluation metrics. Such metrics would not only accelerate the condensation process but also broaden the range of GC applications.
7.4 Robust and Secure Graph Condensation
The condensation of graphs not only facilitates efficient training of GNNs but also contributes significantly to reduce the storage overhead and accelerate the data transmission. However, the robustness of highly compressed graph datasets to various security and privacy threats Sun et al. (2022) (e.g., adversarial attacks, inference breaches, and backdoor attacks) remains unknown. These threats pose significant limitations on the deployment of GC in security-sensitive sectors, including but not limited to confidential communication networks, financial systems and industrial control systems. In light of these challenges, the development of robust, privacy-preserving GC algorithms becomes imperative.
7.5 Explainable Graph Condensation
Another promising avenue for GC research is the development of explainable GC methods. Current GC techniques, while efficient in reducing graph size, often operate as a black box, offering limited insight into how condensed graphs represent the original graph data Yuan et al. (2022). An explainable condensation can illuminate the intricate patterns within the graph, thereby enhancing user trust and acceptance. This is particularly important in sectors like healthcare Wang et al. (2020) and finance Wang et al. (2021), where the interpretability and justification of decision-making processes are critical. Consequently, by elucidating the logic behind data transformations, explainable GC has the potential to open up new avenues for its application, establishing GC as a more reliable and transparent tool in data-centric technology fields.
8 Conclusion
This survey presents an up-to-date and comprehensive overview of GC and classifies the existing research into four distinct categories aligned with the proposed essential GC evaluation criteria. It further explores two critical components of GC: optimization strategies and condensed graph generation, facilitating a deep understanding of GC techniques. Additionally, it highlights the current challenges and emerging perspectives in GC, with the objective of inspiring and guiding future research in this evolving field. Overall, this survey seeks to serve as a beacon for researchers and practitioners, guiding them through the evolving landscape of GC and fostering advancements that could reshape the use of graph data in various practical applications.